1) Balance the following electrochemical equation for an acidic environment. CrO4+Zn Zn2 +Cr
Answer #1

Given reaction is croda + zn a nta rc&ts (acidic medium) In crowdi cr vridation number is c8+46-) = -2 = CH = 8-2 = +6. ta +Step 6: Balance the elections. Add es, electron deficient side zn zota tał o cxo2 outuse crona +84 +3é c&t? +440 ot 3 ut ste